10th term of the progression -4-1+2+5+--------------

Grade:

5 Answers

Godfrey Classic Prince
11 years ago

Dear jayku kumar,

The Formula for finding the nth term in Arithmetic Progression is given by

a= a1 + (n-1)d

where,

ais the value of the nth term

ais the value of the first term

n is the nth term

d is the difference 

 

In this case, a= -4, d = (a-a1) = 3, n=10

Therefore substituting these values,we get the value of the nth term(an)

Therefore, a= -4 + (10-1)3

                   = -4 +(9)3

                   = -4+ 27

                   = 23

Hence the 10th term is 23.

 

 

Hope this helped you immensely...!!

All the Very Best & Good Luck to you ...!!

 

Best Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-M

Please approve my answer if you liked it by clicking on "Yes" given below..!!Smile


 

 

 

Utkarsh Verma
11 years ago

Here, a=-4

and, d = -1-(-4) = 3

a10 = a+(n-1)d = -4+(10-1)(3) = -4+27 = 23.

Therefore, the 10th term is 23.

mahesh dasari
11 years ago

It is in A.P then a = -4 , d=3 then 10th term is  a+( n - 1 ) X d ,   - 4 + 9X 3 = -4 + 27 = 23

shubham chhabra
11 years ago

T=A+(n-1)d

A= -4 , n=10 , d= -1-(-4)=3

T=-4+(10-1)3= 23

Achyut Srivastava
11 years ago

23...

a=-4,

n=10,

d=-1+4=+3,

a=l+(n-1)d..therefore l=-4+27=23